OpenStudy (anonymous):

a rectangle has length triple the side of a square and width two units less than the side of the square. if the two areas are equal, what are the square's dimensions?

OpenStudy (anonymous):

the area of the square is x^2 The rectangle has an area of : 3x * (x-2) The two areas are the same so: x^2=3x^2-6x Thus: -2x^2=-6x Divide both sides by -2 and get: x^2=3x, now divide both sides by x and get: x=3, thus the square has sides of lenght 3.

OpenStudy (anonymous):

thank you very much!